The electric motor of a model train accelerates the train from rest to 0.500 m/s in 28.0 ms. The total mass of the train is 650 g. Find the average power delivered to the train during its acceleration.
Answer #1

answer) here the formula for power is

P=E/t

E=1/2mv2=1/2*0.650*0.5002=0.08125 J

t=28ms=0.028s

so using power formula

P=0.08125/0.028 s =2.90 W

so the answer is 2.90 W or 2.9 W or 3.0 W